# Local Thermal Operations and Classical Communication

### v1

@@ Abstract
Quantum mechanics and thermodynamics are foundational to modern physics. This paper introduces a novel framework, based on local operations with classical communication, fundamental paradigm in quantum information, incorporating the preservation of thermal equilibrium: local thermal operations and classical communication (LTOCC). We introduce a hierarchy of LTOCC, ordered according to the amount of control and communication. Next, we provide a detailed analysis of symmetric local thermal operations and classical communication ($\mathcal{S}$LTOCC), including majorization-based bounds on possible transitions and properties such as generation of strong correlations from Gibbs states.
@@ 1 Introduction
Amongst the pillars on which modern physics stands, quantum mechanics and thermodynamics stand as two of the most robust ones. As sir Arthur Eddington has famously stated, “(…) if your theory is found to be against the second law of thermodynamics I can give you no hope; there is nothing for it but to collapse in deepest humiliation [1].” Similarly, despite significant controversies regarding its ontological status, quantum mechanics has withstood all experimental inquiry it has been subjected to over the century of its progress.
@@
Recent years have seen a growth of the efforts focused on joining these two fields in what is called the quantum thermodynamics [2, 3, 4], with results applicable to the problems of quantum battery charging [5, 6, 7, 8, 9, 10, 11, 12, 13], cooling [14, 15, 16, 17, 18, 19, 20], autonomous quantum machines [21, 22, 23, 24] and more. Most importantly in the current context, however, are efforts to provide ultimate bounds on transformations of low-dimensional systems in contact with heat bath that would respect the conservation of thermal equilibrium. This approach has resulted in introducing the resource-theoretic framework of thermal operations, with wide body of related works already present within the literature [25, 26, 27, 28, 29, 30, 31, 32]
@@
In parallel, the discovery of quantum entanglement and the related notion of quantum nonlocality lead to realization of the fundamental difference between classical and quantum distributed system, with diverse applications, such as teleportation [33], quantum key distribution [34], quantum secret sharing [35] and quantum network nonlocality [36], to name but a few. Early on it has been also realised, that local operations and classical communication (LOCC) is set of transformations, that effected on a nonlocal state do not create entanglement. Thus, a resource theory of entanglement based on LOCC as the set of free operations has been developed and studied in depth, leading to results on entanglement distillation, conversion and others.
@@
In this work, following sir Eddington’s advice, we propose a framework for operations limited both by locality and communication restricted to classical means, as well as incorporating thermodynamic laws, thus adhering to the basic requirements of thermodynamics. For simplicity, we will refer to such operations as local thermal operations and classical communication (LTOCC). In Section 2 we recall all the notions necessary to introduce our results, including noisy and thermal operations and cooling maps, local operations and classical communication, and concepts connected to stochastic tensors and multistochasticity. In Section 3 we introduce the proper framework of LTOCC together with a hierarchy of processes requiring different amounts of communication and control. Finally, in Section 4 we focus on a detailed analysis of a restricted scheme, with additional symmetricity constraints symmetric local thermal operations and classical comminication ($\mathcal{S}$LTOCC). We discuss the problem of extremal channels, consider possible transitions in a fashion similar to the prior results for thermal operations and examine the unexpected properties of $\mathcal{S}$LTOCC such as the creation of strong correlations even for Gibbs states as inputs. Section 5 serves as a brief summary of the proposed framework and sets up a list of open problems worth further investigation.
@@
Figure 1: Artists depiction of Alice and Bob using classical communication, in contact with local thermal baths, thus realising the backbone of LTOCC, Local Thermal Operations and Classical Communication (Courtesy of A. de Oliveira).
@@ 2.1 Noisy operations, thermal operations and cooling maps
Let us begin by introducing the essential frameworks that describe evolution of finite-dimensional quantum states under thermodynamical constraints coming from the contact with a thermal bath characterised by inverse temperature $\beta=(k_{B}T)^{-1}$. The primary system of dimension $d$, prepared in an initial state $\rho$ is characterised by its Hamiltonian $H=\sum_{i}E_{i}\outerproduct{E_{i}}{E_{i}}$ |, where we assume the energy levels $E_{i}$ to be non-degenerate. The total system-bath initial state is assumed to be uncorrelated, $\rho\otimes\gamma_{E}$, with the bath characterised by the Hamiltonian $H_{E}$ starting in the Gibbs state defined as
@@
Moreover, to ensure energy conservation, it is assumed that the unitary operation $U$ commutes with the joint Hamiltonian,
@@
In order to provide proper background to the main topic of this manuscript we need to consider the conditions for the convertibility of states. In the case of finite temperatures, given a pair of states $\rho$ and $\sigma$, the transition is possible if and only if there exists a thermal operation $\mathcal{E}$ such that $\mathcal{E}(\rho)=\sigma$ and $\mathcal{E}(\gamma)=\gamma$. However, as we limit ourselves to energy-incoherent states, the condition reduces to existence of a stochastic matrix $\Lambda$ acting on the respective populations $\mathbf{p}=\operatorname{diag}_{H}(\rho),\,\mathbf{q}=\operatorname{diag}_{H}(\sigma)$ such that
@@
In the particular case of infinite temperature, $\beta=0$, the Gibbs state reduces to the maximally mixed state, with constant populations $\eta_{i}=\frac{1}{d}$, and the Gibbs-preserving condition $\Lambda\boldsymbol{\eta}=\boldsymbol{\eta}$ defines the set of noisy operations acting on the energy-incoherent states equivalent to the set of bistochastic matrices. The state convertibility is then governed by majorization relation.
@@
It may happen that given two vectors, $\mathbf{p}$ and $\mathbf{q}$, are incomparable, meaning that $f^{\,\beta}_{\mathbf{p}}$ and $f^{\,\beta}_{\mathbf{q}}$ cross at some point (see Fig. 2). Furthermore, the thermomajorisation order cannot be seen as a lattice in the sense that for $\mathbf{p}$ and $\mathbf{q}$, which do not share the same $\beta$-order, there is no unique join or meet [40].
Figure 2: Example of majorization and thermomajorization curves for a pair of distributions $\mathbf{p}=(0.32,0.45,0.02,0.21),\,\mathbf{q}=(0.25,0.4,0.2,0.15)$ in $d=4$ with energy levels $E_{i}=i$. In both examples, the distribution $\mathbf{p}$ corresponding to the blue curve majorizes the distribution $\mathbf{q}$ corresponding to the orange curve, $\mathbf{p}\succ\mathbf{q},\,\mathbf{p}\succ_{\beta}\mathbf{q}$, since the blue curve lies entirely above the orange one.
We can now state the generalisation of Theorem 1 from bistochastic matrices to Gibbs-preserving matrices which provides the necessary and sufficient conditions for the existence of a thermal operation between two energy-incoherent states [27, 41].
@@
Bistochastic matrices correspond to thermal operations in infinite temperature. However one can also study the opposite zero temperature limit. In the A we summarize the concept of cooling maps, originating for such a study [42], in a scenario with diagonal quantum states. Furthermore in B we show an alternative derivation of cooling maps for diagonal quantum states as a zero temperature limit of Gibbs-preserving matrices.
In further discussion, following the study of thermal operations, we restrict ourselves to the states that are diagonal in the Hamiltonian eigenbasis. This lets us identify quantum states with their populations $\rho=\sum_{i}p_{i}\;|E_{i}\rangle\langle E_{i}|\Longleftrightarrow\bf{p}$.
@@ 2.2 Local operations and classical communication
Whenever considering multipartite quantum system composed of spatially separated subsystems, each with local Hilbert space $\mathcal{H}_{i}$, it is natural to limit the operations that can be effected on the total system. Such scenario is often described as "distant laboratories" paradigm. The most rigorous restriction comes from allowing the involved parties to perform operations only locally. Thus, for a bipartite system $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ the allowed operations would be limited to channels of the form $\mathcal{E}_{A}\otimes\mathcal{E}_{B}$.
It is however natural to note that distant laboratories do not necessarily mean disconnected laboratories, which leads to inclusion of classical communication, such as shared randomness or exchange of results of experiments. Together with local operations, possibly conditioned on the results of the local measurements, is referred to as local operations and classical communication (LOCC) [43]. Some of the most prominent protocols in quantum information, including quantum teleportation [44] and superdense coding [45], rely on LOCC operations; furthermore, the study of convertibility of entangled states under LOCC, which are known not to increase entanglement, is one of the cornerstones of the quantum resource theories [46, 47].
Despite their widespread use and intuitive understanding, the structure of the set of LOCC operations is intricate – we refer the reader interested in detailed description to [43], where LOCC is introduced with full rigorousness within the framework of quantum instruments. In this work, however, the most important aspect of LOCC is possibility of performing local operations conditioned on results of measurement in a distant laboratory.
@@ 2.3 Multistochastic tensors and multi stochastic Birkhoff polytope
Another important foundation for our work are stochastic tensors. Similarly, as stochastic matrix $\Lambda_{ij}$ encompasses the mapping of probability distributions, the stochastic tensor $T_{ijk}$ describe the mapping of two distributions into one,
@@
and the concept can be extended for more distributions by considering tensors of higher rank. To ensure, that the operation is well-defined, all elements of the stochastic tensor must be positive, $T_{ijk}\geq 0$, and sums of its hypercolumns must be equal one, $\sum_{i}T_{ijk}=1$, for any $j,k$. Typically, physically interesting phenomena are described using restricted scenarios such as bistochasticity, which appears naturally in noisy operations or as transition probabilities arising from unitary evolution of quantum systems [48, 49] . In the case of tensors instead of one additional condition for normalisation of rows (hypercolumns in relation to a selected index), one can demand similar conditions to be satisfied by all hypercolumns, which gives rise to multistochastic tensors.
@@
The resemblance between bistochastic matrices and multistochastic tensors suggests, that many of the properties of the former should translate to the latter. However, such a connection is not entirely straightforward and becomes even more troublesome when considering quantum counterpart for both of them [50].
For example, sets of both objects are convex, forming the so-called Birkhoff polytopes [51]. All extremal bistochastic matrices, vertices of Birkhoff polytope, are permutation matrices [52] – all its entries are either zero or one. Thus one would expect, that permutation tensors, in which all entries are equal to either zero or one, also fully describe the extremal tensors. However, as proven in [53], there are also other extremal multistochastic tensors. Note, that while layers of permutation tensors are extremal bistochastic matrices, it no longer holds for other extremal multistochastic tensors.
@@
$T=(T_{0,j,k},T_{1,j,k},T_{2,j,k})=\left(\begin{matrix}1&0&0\\ 0&1&0\\ 0&0&1\\ \end{matrix}\right.\left|\begin{matrix}0&1&0\\ 0&0&1\\ 1&0&0\\ \end{matrix}\right.\left|\begin{matrix}0&0&1\\ 1&0&0\\ 0&1&0\\ \end{matrix}\right)\,,$ (11)
@@
where the reader should imagine the square sub-matrices arranged in a $3\times 3\times 3$ cube, while the example of other extremal tensors is presented below
$T=(T_{0,j,k},T_{1,j,k},T_{2,j,k})=\left(\begin{matrix}0&\frac{1}{2}&\frac{1}{2}\\ \frac{1}{2}&\frac{1}{2}&0\\ \frac{1}{2}&0&\frac{1}{2}\\ \end{matrix}\right.\left|\begin{matrix}\frac{1}{2}&\frac{1}{2}&0\\ 0&\frac{1}{2}&\frac{1}{2}\\ \frac{1}{2}&0&\frac{1}{2}\\ \end{matrix}\right.\left|\begin{matrix}\frac{1}{2}&0&\frac{1}{2}\\ \frac{1}{2}&0&\frac{1}{2}\\ 0&1&0\\ \end{matrix}\right)\,.$ (12)

### v2

@@ Abstract
In quantum thermodynamics, understanding the interplay between locality, thermal constraints, and communication remains an open challenge. In this manuscript, we introduce Local Thermal Operations and Classical Communication (LTOCC), a novel operational framework that unifies the distant laboratories paradigm with thermodynamic restrictions, defining the fundamental limits on transformations between spatially separated systems. We establish a hierarchy of LTOCC protocols, demonstrating inclusion relations between different levels and revealing their deep connection to semilocal thermal operations. To formalize this framework, we develop thermal tensors and bithermal tensors, extending tristochastic tensors to thermodynamic settings and providing new mathematical tools for constrained quantum processes. Furthermore, we explore the role of LTOCC in approximating logical gates such as CNOT and SWAP and investigate its potential to generate correlations between distant systems.
: New J. Phys.
@@ 1 Introduction
Amongst the pillars on which modern physics stands, quantum mechanics and thermodynamics stand as two of the most robust ones. As Sir Arthur Eddington has famously stated, “(…) if your theory is found to be against the second law of thermodynamics, I can give you no hope; there is nothing for it but to collapse in deepest humiliation [1].” Similarly, despite significant controversies regarding its ontological status, quantum mechanics has withstood all experimental inquiry it has been subject to over the century of its progress.
@@
Recent years have seen a growth of the efforts focused on joining these two fields in what is called quantum thermodynamics [2, 3, 4], with results applicable to the problems of quantum battery charging [5, 6, 7, 8, 9, 10, 11, 12, 13], cooling [14, 15, 16, 17, 18, 19, 20], autonomous quantum machines [21, 22, 23, 24] and more. Most important in the current context, however, are efforts to provide ultimate bounds on transformations of low-dimensional systems in contact with a heat bath that would respect the conservation of thermal equilibrium. This approach has resulted in the introduction of the resource-theoretic framework of thermal operations, with a wide body of related works already present within the literature [25, 26, 27, 28, 29, 30, 31, 32], extended by a concept of semilocal thermal operations, where local thermal baths in different temperatures interact only via local proxy systems, but interactions between the said systems can be arbitrarily nonlocal, classical or quantum, within the thermodynamical limits [33, 34].
@@
In parallel, the discovery of quantum entanglement and the related notion of quantum nonlocality lead to the realization of the fundamental difference between classical and quantum distributed systems, with diverse applications, such as quantum teleportation [35], quantum key distribution [36], quantum secret sharing [37] and quantum network nonlocality [38], to name but a few. Early on it has also been realised that local operations and classical communication (LOCC) constitute a class of transformations that, when applied on a nonlocal state, do not create entanglement.
@@
In this work, following Sir Eddington’s advice, we propose a framework for operations limited both by locality and communication restricted to classical means, as well as incorporating thermodynamic laws locally for all involved parties, thus adhering to the basic requirements of thermodynamics. For simplicity, we will refer to such operations as local thermal operations and classical communication (LTOCC). In Section 2 we recall all the notions necessary to introduce our framework, including: noisy and thermal operations and cooling maps, local operations and classical communication, and concepts connected to stochastic tensors and multistochasticity. Furthermore, in Section 3 we discuss the concept of semilocal thermal operations (SLTO) and present new results, which are required to establish a connection between SLTO and LTOCC. In Section 4 we introduce the proper framework of LTOCC together with a hierarchy of protocols requiring different amounts of communication and control, demonstrating their relationship with semilocal thermal operations. In Section 5 we study the basic building blocks of the framework – thermal tensors – with a special focus on the nontrivial, symmetric subset of bithermal tensors, which are thermal counterparts of tristochastic tensors. Finally, in Section 6 we present thermal counterparts of standard logical operations which fall within the introduced framework and discuss the correlating abilities of one-round LTOCC and symmetric LTOCC. Section 7 serves as a brief summary of the proposed framework and sets up a list of open problems for future investigation. A and B present a discussion of cooling maps and their alternative derivation, C provides a detailed discussion of one- and two-round LTOCC protocols with and without memory, and D presents new, auxiliary results on semilocal thermal operations and proof of inclusion relation between LTOCC and SLTO.
@@
Figure 1: Artist’s depiction of Alice and Bob using classical communication, in contact with local thermal baths, thus realising the backbone of LTOCC, Local Thermal Operations and Classical Communication (Courtesy of A. de Oliveira Junior).
Summary of results
Due to the extent of this manuscript, we provide a concise summary of the novel results provided:
  1. 1.
    Proofs of convexity and closedness under composition for semilocal thermal operations (SLTO) (Theorems 4 and 5)
  2. 2.
    Introduction of a new paradigm of local thermal operations with classical communication (LTOCC) and corresponding communication complexity hierarchy, together with proofs of inclusion between different strata of hierarchy (Section 4.3)
  3. 3.
    Proof of inclusion of LTOCC with shared randomness in the SLTO in a general quantum scenario for any number of rounds (Theorem 7) and conjecture concerning equality of the sets in the infinite-round limit for energy-incoherent states (Conjecture 1)
  4. 4.
    Detailed study of basic building blocks of the framework – thermal and bithermal tensors – with particular attention to the latter as analogues of tristochastic tensors, with preliminary results on the structure of thermal analogue of tristochastic Birkhoff polytope (Section 5)
  5. 5.
    Case studies of the use of no-memory LTOCC to approximate CNOT and SWAP logic gates (Section 6.1) and the possibility of creating significant correlations between the involved systems when using memory-extended LTOCC (Section 6.2)
Additionally, in C we present a detailed consideration of 1- and 2-round LTOCC together with placing them in the context of sets of free states for composite resource theories as defined in [39].
@@ 2.1 Noisy operations, thermal operations and cooling maps
Let us begin by introducing the essential frameworks that describe the evolution of finite-dimensional quantum states under thermodynamical constraints resulting from contact with a thermal bath, characterised by inverse temperature $\beta=(k_{B}T)^{-1}$ which will also govern local evolution of our subsystem. The primary system of dimension $d$, prepared in an initial state $\rho$ is characterised by its Hamiltonian $H=\sum_{i}E_{i}\outerproduct{E_{i}}{E_{i}}$ |, where we assume the energy levels $E_{i}$ to be non-degenerate. The total system-bath initial state is assumed to be uncorrelated, $\rho\otimes\gamma_{E}$, with the bath characterised by the Hamiltonian $H_{E}$ starting in the Gibbs state defined as
@@
Moreover, to ensure total energy conservation, it is assumed that the unitary operation $U$ commutes with the joint Hamiltonian,
@@
[…]
@@
In the following paragraphs, we discuss the resulting restrictions on possible transitions between energy-incoherent states, $\mathbf{p}\rightarrow\mathbf{q}$. Note that the following considerations apply as necesery conditions also whenever the original state $\rho$ is coherent and $\operatorname{diag}(\rho)=\mathbf{p}$ [27].
In the particular case of infinite temperature (noisy operations), $\beta=0$, the Gibbs state reduces to the maximally mixed state, with constant populations $\eta_{i}=\frac{1}{d}$, and the Gibbs-preserving condition $\Lambda\boldsymbol{\eta}=\boldsymbol{\eta}$ defines the set of noisy operations acting on the energy-incoherent states equivalent to the set of bistochastic matrices. The state convertibility is then governed by majorization relation.
@@
It may happen that a given pair of vectors, $\mathbf{p}$ and $\mathbf{q}$, are incomparable, meaning that $f^{\,\beta}_{\mathbf{p}}$ and $f^{\,\beta}_{\mathbf{q}}$ cross at some point (see Fig. 2). Furthermore, the thermomajorisation order cannot be seen as a lattice in the sense that for $\mathbf{p}$ and $\mathbf{q}$, which do not share the same $\beta$-order, there is no unique join or meet, corresponding to the unique first point in the joint future and last point in the joint past [43].
Figure 2: Example of majorization and thermomajorization curves for a pair of distributions $\mathbf{p}=(0.2,0.33,0.3,0.16),\,\mathbf{q}=(0.45,0.09,0.45,0.01)$ in $d=4$ with energy levels $E_{i}=i$ as defined in eq (7). In both examples, the distribution $\mathbf{p}$ corresponding to the blue curve majorizes the distribution $\mathbf{q}$ corresponding to the orange curve, $\mathbf{p}\succ\mathbf{q},\,\mathbf{p}\succ_{\beta}\mathbf{q}$, since the blue curve lies entirely above the orange one. Notice that in infinite temperature $\beta=0$, the thermomajorization problem reduces to majorization one since all elbows of the curves happen for the same values of $x$, so it is sufficient to consider only these points.
We can now state the generalisation of Theorem 1 from bistochastic matrices to Gibbs-preserving stochastic matrices which provides the necessary and sufficient conditions for the existence of a thermal operation between two energy-incoherent states [27, 44].
@@
Bistochastic matrices correspond to thermal operations on energy-incoherent states in infinite temperature. However one can also study the opposite zero temperature limit. In the A we summarize the concept of cooling maps, originating for such a study [45], in a scenario with diagonal quantum states. Furthermore in B we show an alternative derivation of cooling maps for diagonal quantum states as a zero temperature limit of Gibbs-preserving matrices.
In further discussion, following the study of thermal operations, we restrict ourselves to the states that are diagonal in the Hamiltonian eigenbasis, unless stated otherwise. This lets us identify quantum states with their populations $\rho=\sum_{i}p_{i}\;|E_{i}\rangle\langle E_{i}|\Longleftrightarrow\bf{p}$.
@@ 2.2 Local operations and classical communication
Whenever considering a multipartite quantum system composed of spatially separated subsystems, each with local Hilbert space $\mathcal{H}_{i}$, it is natural to limit the operations that can be effected on the total system. Such a scenario is often described as a “distant laboratories” paradigm. The most rigorous restriction comes from allowing the involved parties to perform operations only locally. Thus, for a bipartite system $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ the allowed operations would be limited to channels of the form $\mathcal{E}_{A}\otimes\mathcal{E}_{B}$.
It is, however, natural to note that distant laboratories should not be necessarily equated to disconnected laboratories, which leads to inclusion of classical communication, such as shared randomness or exchange of results of experiments. Together with local operations, possibly conditioned on the measurement results, it is referred to as local operations and classical communication (LOCC) [46]. Some of the most prominent protocols in quantum information, including quantum teleportation [47] and superdense coding [48], rely on LOCC operations; furthermore, the study of convertibility of entangled states under LOCC, which are known not to increase entanglement, is one of the cornerstones of the quantum resource theories [49, 50].
Despite their widespread use and intuitive understanding, the structure of the set of LOCC operations is intricate – we refer the reader interested in detailed description to [46], where LOCC is introduced with full rigorousness within the framework of quantum instruments. In this work, however, the most important aspect of LOCC is a possibility of performing local operations conditioned on results of measurement in a distant laboratory.
@@ 2.3 Multistochastic tensors and multi stochastic Birkhoff polytope
Another important foundation for our work are stochastic tensors. Similarly, as stochastic matrix $\Lambda_{ij}$ encompasses the mapping of probability distributions, a stochastic tensor $T_{ijk}$ describes the mapping of two distributions into one,
@@
and the concept can be extended for more distributions by considering tensors of higher rank. To ensure, that the operation is well-defined, all elements of the stochastic tensor must be positive, $T_{ijk}\geq 0$, and sums of its hypercolumns are normalised, $\sum_{i}T_{ijk}=1$, for any $j,k$. Typically, physically interesting phenomena are described using restricted scenarios such as bistochasticity, which appears naturally in noisy operations or as transition probabilities arising from unitary evolution of quantum systems [51, 52]. In the case of tensors instead of one additional condition for the normalisation of rows (hypercolumns in relation to a selected index), one can demand similar conditions to be satisfied by all hypercolumns, which gives rise to multistochastic tensors.
@@
The resemblance between bistochastic matrices and multistochastic tensors suggests, that many of the properties of the former should translate to the latter. However, such a connection is not entirely straightforward and becomes even more troublesome when considering quantum counterparts for both of them [53].
For example, sets of both bistochastic matrices and tristochast tensors are convex, forming the so-called Birkhoff polytopes [54]. All extremal bistochastic matrices, vertices of Birkhoff polytope, are permutation matrices [55] – all its entries are either one or zero. Thus one would expect, that permutation tensors, in which all entries are equal to either one or zero, also fully describe the extremal tensors. However, as proven in [56], there are also other extremal multistochastic tensors that do not correspond to permutation tensors. Note, that while layers of permutation tensors are extremal bistochastic matrices, it no longer holds for other extremal multistochastic tensors.
@@
$T=(T_{0jk},T_{1jk},T_{2jk})=\left(\begin{matrix}1&0&0\\ 0&1&0\\ 0&0&1\\ \end{matrix}\right.\left|\begin{matrix}0&1&0\\ 0&0&1\\ 1&0&0\\ \end{matrix}\right.\left|\begin{matrix}0&0&1\\ 1&0&0\\ 0&1&0\\ \end{matrix}\right)\,,$ (11)
@@
where the reader should imagine the square sub-matrices arranged in a $3\times 3\times 3$ cube, while an example of a non-permutation extremal tensor is presented below
$T=(T_{0jk},T_{1jk},T_{2jk})=\left(\begin{matrix}0&\frac{1}{2}&\frac{1}{2}\\ \frac{1}{2}&\frac{1}{2}&0\\ \frac{1}{2}&0&\frac{1}{2}\\ \end{matrix}\right.\left|\begin{matrix}\frac{1}{2}&\frac{1}{2}&0\\ 0&\frac{1}{2}&\frac{1}{2}\\ \frac{1}{2}&0&\frac{1}{2}\\ \end{matrix}\right.\left|\begin{matrix}\frac{1}{2}&0&\frac{1}{2}\\ \frac{1}{2}&0&\frac{1}{2}\\ 0&1&0\\ \end{matrix}\right)\,.$ (12)
3 Interlude: Auxiliary results on semilocal thermal operations
Recently, there has been an effort to formulate a notion similar to thermal operations, which would be applicable to distributed systems interacting with different local thermal baths [33]. This approach resulted in the construction of the theory named semilocal thermal operations (SLTO), which is suitable to consider the ultimate limitations of finite-size thermal engines. This framework is very general, incorporating simultaneous interactions between all system and bath elements.
In our work we treat SLTO as an important background, which operates within minimal assumptions similar to thermal operations, but does not provide any additional structural restrictions on the manner in which the involved systems interact. The exact definition of semilocal thermal operation is as follows:
Definition 4 ([33] Definiton 1).
Let us consider a system $S$ consisting of two subsystems $S^{(A)}$, $S^{(B)}$ with Hamiltonian $H=H^{(A)}$ + $H^{(B)}$. An operation $\Lambda^{(AB)}$ with respect to local inverse temperatures $\beta^{(A)}$ and $\beta^{(B)}$ is called semilocal thermal operation if there exist two local thermal baths $\mathcal{B}^{(A)}$, $\mathcal{B}^{(B)}$ with Hamiltonians $H^{\mathcal{B}^{(A)}}$, $H^{\mathcal{B}^{(B)}}$ such that
$\Lambda^{(AB)}(\rho^{(AB)})=\text{Tr}_{\mathcal{B}^{(A)},\mathcal{B}^{(B)}}\left[U(\gamma^{\mathcal{B}^{(A)}}\otimes\gamma^{\mathcal{B}^{(A)}}\otimes\rho^{(AB)})U^{\dagger}\right]$ (13)
where the global unitary matrix $U$ satisfies the following commutation relations:
$[U,H^{(A)}+H^{(B)}+H^{\mathcal{B}^{(A)}}+H^{\mathcal{B}^{(B)}}]=0,$ (14a)
$[U,\beta^{(A)}(H^{(A)}+H^{\mathcal{B}^{(A)}})+\beta^{(B)}(H^{(B)}+H^{\mathcal{B}^{(B)}})]=0$ (14b)
The first constraint (14a) imposes conservation of total system-bath energy across both systems. To understand the second one, one should consider the systems and baths in the thermodynamic limit. Then, the energy flow from or to the thermal bath corresponds to heat exchange, which in this limit is equal to entropy change times bath’s temperature. Thus, the second equation (14b), conservation of weighted heat flow, is in fact encoding the “second law of thermodynamics” – entropy conservation. We stress that when both thermal baths have the same temperature, semilocal thermal operations are equivalent to thermal operations on a joint system with one joint bath.
The authors of [33] provided a strong operational bound on possible transformations of bipartite states using SLTO:
Theorem 3 ([33], Supplementary Information, Theorem 11).
The transition between two energy-incoherent states $\rho^{(AB)}\to\sigma^{(AB)}$ can occur under semi-local thermal operation if, and only if, the diagonal of $\rho^{(AB)}$ thermo-majorizes the diagonal of $\sigma^{(AB)}$ with respect to state $\gamma^{(A)}\otimes\gamma^{(B)}$
The above theorem can be extended by an important corollary. Let us perform an atrificial rescaling of Hamiltonian $\tilde{H}^{(A)}=\sqrt{\frac{\beta^{(A)}}{\beta^{(B)}}}H^{(B)},\,\tilde{H}^{(B)}=\sqrt{\frac{\beta^{(B)}}{\beta^{(A)}}}H^{(B)}$, so that the Gibbs states with common inverse temperature $\tilde{\beta}=\sqrt{\beta^{(A)}\beta^{(B)}}$ and Hamiltonians $\tilde{H}^{(A)}$ and $\tilde{H}^{(B)}$ are the same as the original Gibbs states $\gamma^{(A)}$ and $\gamma^{(B)}$. Then the condition from Theorem 3, by the Theorem 2, is equivalent to the preservation of global Gibbs state $\gamma^{(AB)}=\gamma^{(A)}\otimes\gamma^{(B)}$ at a common temperature $\tilde{\beta}=\sqrt{\beta^{(A)}\beta^{(B)}}$. Thus, we have.
Corollary 1 ([33], Supplementary Information, Corollary 4).
The transition between two energy-incoherent states $\rho^{(AB)}\to\sigma^{(AB)}$ can occur under semi-local thermal operation if, and only if, there exist stochastic matrix preserving product of Gibbs states $M(\gamma^{(A)}\otimes\gamma^{(B)})=\gamma^{(A)}\otimes\gamma^{(B)}$, that maps the spectrum of one state into the other $M(\text{diag}(\rho^{(AB)}))=\text{diag}(\sigma^{(AB)})$.
Therefore, if one discusses only energy-incoherent states, SLTO can be effectively treated as stochastic Gibbs-preserving matrices. Note that those matrices are defined by simple linear equations from Corollary 1 and stochasticity conditions, which let us state new corollary:
Corollary 2 (Set-closure of SLTO).
The set of possible transitions between finite-dimensional energy-incoherent states $\rho^{(AB)}\to\sigma^{(AB)}$ under SLTO is closed.
We present two more properties of SLTO, which are crucial from the perspective of resource theories and important in the future work but, to our knowledge, were not proven before as well.
Theorem 4.
(Composition-closure of SLTO) The set of semilocal thermal operations is closed under composition, i.e if $\Lambda_{1}^{(AB)}$ and $\Lambda_{2}^{(AB)}$ are SLTO then $\Lambda_{2}^{(AB)}\circ\Lambda_{1}^{(AB)}$ is SLTO with the same pair of inverse temperatures $\beta^{(A)},\,\beta^{(B)}$.
Theorem 5.
(Convexity of SLTO) The set of semilocal thermal operations is convex, i.e if $\Lambda_{1}^{(AB)}$ and $\Lambda_{2}^{(AB)}$ are SLTO then $\alpha\Lambda_{1}^{(AB)}+(1-\alpha)\Lambda_{2}^{(AB)}$ is SLTO with the same inverse temperatures for any $\alpha\in[0,1]$.
Proofs of Theorems 4 and 5 are provided in D. ∎
In what follows we will establish relation between the general framework of SLTO and a communication-constrained scenario of local thermal operations and classical communication, which will be described at length in the sections to follow.